An aqueous solution of 2% non-volatile exerts a pressure of 1.004 Bar at the normal boiling point of the solvent. What is the molar mass of the solute?
1
Expert's answer
2020-05-14T14:33:00-0400

Here,

Vapour pressure of the solution at normal boiling point "(p_1)=1.004\ bar"

Vapour pressure of pure water at normal boiling point "(p_1^0)=1.013\ bar"

Mass of solute,"(w_2) = 2 \ g"

Mass of solvent (water),"(w_1) = 100 - 2 = 98 \ g"

Molar mass of solvent (water), "(M_1) = 18\ g \ mol ^{- 1}"

According to Raoult's law,

"\frac{(p_1^o - p1)} {p_1^o } = \frac{ (w_2 \times M_1 )} {(M_2 \times w_1 )}" "\implies \frac{(1.013 - 1.004)} {1.013 }= \frac{(2 \times 18)}{(M_2 \times 98 )}" "\implies M_2=41.35\ g\ mol^{-1}"